OpenStudy (anonymous):

Solve for x: −4|x + 5| = −16 x =21/4 , x = −11/4 x = −1, x = 9 x = −1, x = −9 No solution

OpenStudy (anonymous):

i believe its x=-1, x=9

OpenStudy (gorv):

-4|x+5|=-16 |x+5|=4 \[x+5=\pm4\]

OpenStudy (gorv):

x+5=+4 x=-1 and x+5=-4 x=-9